OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment PQ when the coordinates are P(-12,-9) and Q(2,10).

OpenStudy (angela210793):

XM=(x1+x2)/2 and YM=(Y2-Y1)/2 just replace X1=-12 X2=2 Y1=-9 Y2=10 XM=-10/2=-5 YM=1/2 M(-5;1/2)